A bag contains 5 white balls and 6 red balls. Three balls are drawn, without replacement, from the bag. What is the probability that at least one ball is white?

Answer by stanbon(75887)

A bag contains 5 white balls and 6 red balls. Three balls are drawn, without replacement, from the bag. What is the probability that at least one ball is white?
------
Ans: P(>= 1 white) = 1 - P(no white)
----
= 1 - [5C3/11C3]
----
= 1 - 10/165
----
= 155/165
----
= 0.9394
==============
Cheers,
Stan H.
================